OpenStudy (anonymous):

Write the first five terms of the arithmetic sequence. a1=-2, d=5

jimthompson5910 (jim_thompson5910):

first term is -2 (given) the second term is found by adding 5 to the previous term to get -2+5 = -3 the third term is found by adding 5 to the second term: -3+5 = 2 etc etc

jimthompson5910 (jim_thompson5910):

oh sry typo

jimthompson5910 (jim_thompson5910):

second term: -2+5 = 3 third term: 3 + 5 = 8 etc etc

OpenStudy (anonymous):

thank you! i got it now :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (jhannybean):

\[\large a_{n}= a_{1}+(n-1)d\] so in order to figure out the rest of your terms just plug in what n (number) you're looking for with the rest of your givens.

OpenStudy (jhannybean):

\[\large a_{2}= -2 +(2-1)(5)\] etc.

OpenStudy (dan815):